OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y -3x - 2? A. (0, -1) B. (0, 0) C. (-1, 0) D. (1, 1) How would I solve this??

OpenStudy (anonymous):

y > -3x - 2 or y < -3x - 2 ? ?

OpenStudy (anonymous):

y >= -3x - 2.... sorry lol

OpenStudy (anonymous):

for the first one plug in x=0 and y = -1 -1 >= 0 - 2 or -1 >= -2 Is -1 greater than -2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then A is our answer

OpenStudy (anonymous):

oooo okay so I would do the same for this one: Which of the following ordered pairs is a solution to the inequality y >= 3x - 2? A. (2, 0) B. (0, 4) C. (0, -3) D. (1, -1)

OpenStudy (anonymous):

and for the first one it asks which inequality is NOT a solution. so is A still right?